KyledaGreat:

Consider the following relation. y=x−5 Find four points contained in the inverse. Express your values as an integer or simplified fraction.

KyledaGreat:

{( ), ( ),( ),( )}

KyledaGreat:

@tranquility

Tranquility:

Switch the x and y in the equation. And then solve for y For example If you had y = 2x + 4 You switch the x and y and you get: x = 2y + 4 You make the `x` a `y` and the `y` an `x`. Now solve for y. 2y = x - 4 y = x/2 - 4 That's how you find the inverse

Tranquility:

What's the inverse for y = x - 5 ? To find four points on the inverse, you then plug in any four numbers you want and find the resulting value. If that's too hard for you, then you can graph it using https://www.desmos.com/calculator and you can pick any four points off of the graph for the inverse